OpenStudy (anonymous):

There are 526 tickets sold for a basketball game. The activity cardholders tickets cost 2.25 and the non cardholders tickets cost 3.25. the total amount of money collected was $1034.50. How many of each kind of ticket were sold?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

say, # of card holder tickets = \(c\) # of noncard holder tickets = \(n\)

ganeshie8 (ganeshie8):

lets try and translate the given information into system of equations

ganeshie8 (ganeshie8):

`There are 526 tickets sold for a basketball game` \(\large c + n = 526 \tag{1}\)

ganeshie8 (ganeshie8):

`The activity cardholders tickets cost 2.25 and the non cardholders tickets cost 3.25. the total amount of money collected was $1034.50.` \(\large 2.25c + 3.25n = 1034.5 \tag{2}\)

ganeshie8 (ganeshie8):

two equations and two unknowns anybody can solve

OpenStudy (mokeira):

letcardholders e m and non cardholders be n \[m+n=526\] \[2.25m+3.25n=1034.50\] \[m=526-n\] substitute \[2.25(526-n)+3.25n=103.50\] simplify from here

OpenStudy (midhun.madhu1987):